# Propagation and emission of gravitational waves in the weak-field limit within the Palatini formalism

### v1

@@ V Emission of gravitational waves in the weak-field limit
$Q^{kl}=\frac{1}{\phi_{0}}\int d^{3}x^{\prime}\left(T^{00}-\frac{1}{\phi_{0}V_{0}^{\prime\prime}}\partial_{j}\partial^{j}T\right)x^{k\prime}x^{l\prime}$ (40)
is the quadrupole moment for Palatini $f(R)$ gravity. Then, the expression for the gravitational wave (34) at the quadrupole order yields:
@@
$\left[\theta_{ij}^{TT}\right]_{\text{quad}}=\frac{4G}{rc^{4}}\Lambda_{ijkl}\ddot{Q}^{kl}\ .$ (41)
The first term in the r.h.s. of (40) corresponds to the usual one of GR but with an effective gravitational constant given by $G_{eff}=\frac{G}{\phi_{0}}$. Nevertheless, the second term turns out a new contribution that just shown up in this class of gravitational theories. Such contribution is removed for a homogeneous and isotropic energy-momentum tensor, but might become important in other cases. However, even whether the tensor $T^{\mu\nu}$ represents a inhomogeneous and/or anisotropic source, one would expect that variations along different spatial directions might be small enough to become negligible in the weak-field limit. Nevertheless, such a term depends on the vacuum expectation value of the scalar field potential and specifically on its second derivative, that can be defined as the scalar field effective mass:
$m_{eff}^{2}=V^{\prime\prime}(\phi_{0})\ .$ (42)
Hence, one can conclude that for massive scalar fields, the second contribution in (40) can be neglected while light fields might provide an important contribution to the generation of GWs, even whether the source is nearly homogeneous and isotropic but not completely. In addition, at the quadrupole order the energy radiated with the gravitational wave is:
@@ VI Conclusions
Along this paper, the analysis of GWs propagation and emission is carried out for $f(R)$ theories within the Palatini formalism. To do so, firstly the field equations are expressed in a scalar-tensor-like form, which lead to the presence of a non-dynamical scalar field but sourced by the presence of matter, according to the scalar field equation itself. Then, the full set of field equations are linearised at first order in perturbations of the spacetime metric and the scalar field over an arbitrary background. To analyse the weak-field regime, we focus then on studying first order tensor perturbations of the metric in vacuum and also with the presence of weak enough sources. For the former case, the usual wave equation of GR is recovered, such that GWs propagation within the Palatini formalism obey the same rules as in GR with respect to polariations and speed of propagation.
Nevertheless, when assuming the presence of sources and analysing the emission of GWs through retarded Green functions, one finds that the multipoles expansion not only depend on time variations of the energy density, as provided by the 00-component of the energy-momentum tensor but also in spatial derivatives of its trace. Hence, for homogeneous and isotropic sources, the same results as in GR are recovered and the lowest order in the multipole expansion (the quadrupole moment) leads to the dependence as in GR. However, the presence of inhomogeneous sources introduces inevitable corrections to every order at the expansion and particularly, it modifies the quadrupole moment, leading to a stronger or weaker GWS emission, depending on the source. We suggest that for an enough massive scalar field, such corrections are neglected, while light scalar fields are ruled out, since will introduce strong deviations, making the energy radiated by the source too weak or too strong to satisfy observational constraints.
Next steps suggest that GWs must be analysed as well in strong-field regimes within Palatini-like theories, since the analysis of this paper suggest that while in vacuum, usual perturbation equations of GR are recovered, this changes completely with the presence of sources. In addition, the extension of these analysis to arbitrary RBG theories will provide a grateful insight on the understanding of theories within the Palatini formalism.

### v2

@@ V Emission of gravitational waves in the weak-field limit
$Q^{kl}=\frac{1}{\phi_{0}}\int d^{3}x^{\prime}\left(T^{00}-\frac{1}{\kappa^{2}\phi_{0}V_{0}^{\prime\prime}}\partial_{j}\partial^{j}T\right)x^{k\prime}x^{l\prime}$ (40)
is the quadrupole moment for Palatini $f(R)$ gravity. The first term in the r.h.s. of (40) corresponds to the usual quadrupole of GR but with an effective gravitational constant given by $G_{eff}=\frac{G}{\phi_{0}}$ while the second one is an extra contribution that arises in this class of gravitational theories. However, this second term turns out to vanish in the transverse-traceless gauge, when it is substituted in equation (35). Let us show this fact by integrating this term by parts:
$\int d^{3}x^{\prime}\partial_{j}\partial^{j}Tx^{k\prime}x^{l\prime}$ $=$ $\int d^{3}x^{\prime}\partial_{j}\left(\partial^{j}Tx^{k\prime}x^{l\prime}\right)$
$-\int d^{3}x^{\prime}\partial^{j}\left(T\partial_{j}\left(x^{k\prime}x^{l\prime}\right)\right)$ $+$ $\int d^{3}x^{\prime}T\partial_{j}\partial^{j}\left(x^{k\prime}x^{l\prime}\right)\ ,$ (41)
$=$ $2\int d^{3}x^{\prime}T\delta^{kl}\ ,$
where again the boundary terms are zero because the volume of integration is taken over the source. In this way, the second term in the r.h.s. of (40) is proportional to $\delta^{kl}$ and vanishes when computing (35), since $\Lambda_{ijkl}\delta^{kl}=0$. Therefore, although the GWs propagation equation in Palatini $f(R)$ differs from the one predicted in GR, they are equivalent in the transverse-traceless gauge, at least to quadrupolar order in the multipolar expansion.
At this point, one can continue the analysis by computing higher order terms in the multipolar expansion. In fact, by following a similar procedure as in (38), an expression for $\dot{S}^{klm}$ in terms of $\dddot{T}_{\text{eff}}^{00}$ and $\ddot{T}_{\text{eff}}^{k0}$ can be obtained. With this purpose, we use the definition in (36) for the octupolar term and the fact that the energy-momentum tensor is conserved (37). However, in this case due to the cyclic behaviour of the integration by parts, we can obtain the following two implicit equations:
$\dot{S}^{klm}+\dot{S}^{kml}$ $=$ $-\int d^{3}x^{\prime}\ddot{T}_{\text{eff}}^{k0}x^{l\prime}x^{m\prime}\ ,$ (42)
$\dot{S}^{klm}+\dot{S}^{lmk}+\dot{S}^{mkl}$ $=$ $\frac{1}{2}\int d^{3}x^{\prime}\dddot{T}_{\text{eff}}^{00}x^{k\prime}x^{l\prime}x^{m\prime}\ .$ (43)
By combining both equations, (42) and (43), an explicit expression for $\dot{S}^{klm}$ is obtained:
$\dot{S}^{klm}=\frac{1}{2}\int d^{3}x^{\prime}\dddot{T}_{\text{eff}}^{00}x^{k\prime}x^{l\prime}x^{m\prime}+\int d^{3}x^{\prime}\ddot{T}_{\text{eff}}^{m0}x^{l\prime}x^{k\prime}\ .$ (44)
Here, for convenience, let us split the effective energy-momentum tensor into two terms:
$T_{\mu\nu}^{\text{eff}}=T_{\mu\nu}^{\text{GR}}+T_{\mu\nu}^{\text{PA}}\ ,$ (45)
where,
$T_{\mu\nu}^{\text{GR}}$ $=$ $\frac{1}{\phi_{0}}T_{\mu\nu}\ ,$ (46)
$T_{\mu\nu}^{\text{PA}}$ $=$ $-\frac{1}{\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\left(\nabla_{\mu}\nabla_{\nu}T-\eta_{\mu\nu}\Box T\right)\ .$ (47)
Note that equation (46) is the usual energy-momentum tensor of GR except for the effective gravitational constant while equation (47) is the contribution of Palatini $f(R)$ theories. With that in mind, we can also split (44) into its GR-like part and Palatini $f(R)$ contribution. The latter reads:
$\dot{S}_{\text{PA}}^{klm}=\frac{1}{2}\int d^{3}x^{\prime}\dddot{T}_{\text{PA}}^{00}x^{k\prime}x^{l\prime}x^{m\prime}+\int d^{3}x^{\prime}\ddot{T}_{\text{PA}}^{m0}x^{l\prime}x^{k\prime}\ .$ (48)
Then, by inserting equation (47) into (48), it leads to:
$\dot{S}_{\text{PA}}^{klm}$ $=$ $-\frac{1}{\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\frac{\partial^{3}}{\partial t^{3}}\bigg{[}\frac{1}{2}\int d^{3}x^{\prime}\partial_{j}\partial^{j}Tx^{k\prime}x^{l\prime}x^{m\prime}$ (49)
$+$ $\int d^{3}x^{\prime}\partial^{m}Tx^{l\prime}x^{k\prime}\bigg{]}\ .$
Hence, analogously to (41) and keeping in mind that boundary terms vanish, the expression for $\dot{S}^{klm}$ is reduced to:
$\dot{S}_{\text{PA}}^{klm}=-\frac{1}{\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\frac{\partial^{3}}{\partial t^{3}}\int d^{3}x^{\prime}T\delta^{kl}x^{\prime m}\ .$ (50)
Whereas the corresponding contraction with the normal vector $n$ yields:
$n_{m}\dot{S}_{\text{PA}}^{klm}=-\frac{1}{3\kappa^{2}\phi_{0}^{2}V^{\prime\prime}(\phi_{0})}\frac{\partial^{3}}{\partial t^{3}}\int d^{3}x^{\prime}T\delta^{kl}n_{m}x^{\prime m}\ .$ (51)
Recall that the lambda tensor vanishes when contracted with the Kronecker delta, i.e. $\Lambda_{ijkl}\delta^{kl}=0$, this implies that the Palatini $f(R)$ contribution to the octupolar term to equation (35) vanishes as well. Hence, the octupole order turns out the same as in GR:
$\dot{S}^{klm}=\frac{1}{2}\int d^{3}x^{\prime}\dddot{T}^{00}x^{k\prime}x^{l\prime}x^{m\prime}+\int d^{3}x^{\prime}\ddot{T}^{m0}x^{l\prime}x^{k\prime}\ .$ (52)
One might go further in the multipole expansion, but same results will be obtained. Hence, in the weak-field limit, the multipolar expansion in Palatini $f(R)$ gravity coincide exactly with the GR predictions. In this sense, if one just keeps the lowest order of the expansion (quadrupole order), the expression for the gravitational wave (35) reads:
@@
$\left[\theta_{ij}^{TT}\right]_{\text{quad}}=\frac{4G}{rc^{4}}\Lambda_{ijkl}\ddot{Q}^{kl}\ ,$ (53)
where the quadrupole moment (39) reduces to the one of GR:
$Q^{kl}$ $=\frac{1}{\phi_{0}}\int d^{3}x^{\prime}T^{00}x^{k\prime}x^{l\prime}\ .$ (54)
Since the energy radiated by the GWs emission is given by:
@@ VI Conclusions
Along this paper, the analysis of GWs propagation and emission is carried out for $f(R)$ theories within the Palatini formalism. To do so, firstly the field equations are expressed in a scalar-tensor-like form, which lead to the presence of a non-dynamical scalar field but sourced by the presence of matter, according to the scalar field equation itself. Then, the full set of field equations are linearised at first order in perturbations of the spacetime metric and the scalar field over an arbitrary background. To analyse the weak-field regime, we focus on studying first order tensor perturbations of the metric in vacuum and also with the presence of weak enough sources. For the former case, the usual wave equation of GR is recovered, such that GWs propagation within the Palatini formalism obey the same rules as in GR with respect to polarisations and speed of propagation.
When assuming the presence of sources and analysing the emission of GWs through retarded Green functions, despite that one finds that the GWs equation is sourced by an effective energy-momentum tensor that depends on derivatives of the trace of the energy-momentum tensor, the multipole expansion only depends on time variations of the energy density, as provided by the 00-component of the energy-momentum tensor. Hence, the same results as in GR are recovered and the lowest orders in the multipole expansion (the quadrupole and octupole moments) lead to the same dependence as in GR. One expects that higher orders in the multipole expansion will lead behave in the same way. A remarkable point is that these results are independent of the form of the gravitational action, such that any $f(R)$ action within the Palatini formalism will lead to the same multipole expansion as in GR.
Nevertheless, here we have considered just weak sources and assumed a flat background on which the perturbations arise. One would expect that for strong-field regimes, the emission of GWs will differ from GR predictions and will depend somehow on the gravitational action $f(R)$ or in other words, on the potential for the scalar field. The initial expressions for the quadrupole (39) and octupole moments (49) show an inverse dependence with respect to the effective scalar field mass $m_{eff}^{2}=V^{\prime\prime}(\phi_{0})$ that suggest that in strong field regimes, light scalar field might introduce large deviations with respect to GR predictions whereas for an enough massive scalar field, such corrections might be neglected.
Within this framework, next steps suggest that GWs must be analysed as well in strong-field regimes for Palatini-like theories, since the analysis of this paper suggest that while in linear theory, usual perturbation equations of GR are recovered, this might change completely with the presence of strong enough sources. In addition, the extension of these analysis to arbitrary RBG theories will provide a grateful insight on the understanding of theories within the Palatini formalism.